OpenStudy (anonymous):

add. simplify if possible x + 1 ---- ---- x^2-2x+ x^2-5x+4

OpenStudy (anonymous):

If I assume there is a 1 missing from the end of your first denominator, then \[\dfrac{x}{(x-1)^2} +\dfrac{1}{(x-4)(x-1)}=\dfrac{1}{(x-1)} \bigg( x+ \dfrac{1}{(x-4)} \bigg)\] \[=\dfrac{1}{(x-1)} \bigg( \dfrac{x^2-4x+1}{(x-4)} \bigg)= \dfrac{x^2-4x+1}{(x-1)(x-4)}\] which cannot be simplified further. (Heh, I have just realized the LaTeX works, sortah)